# Conformal Prediction Adaptive to Unknown Subpopulation Shifts

### v1

@@ 6.2 Results for vision tasks
Coverage with varying test environments.
We obtain the results for Algorithm 3 with the same setup as section 5.2 and the results are plotted in Figure  3. From Figure  3 (Left) we observe that Algorithm 3 was able to provide coverage for the majority of test environments while standard conformal prediction could not for a significant number of test environments. Although not as small as Algorithm 1 and 2, Algorithm 3 is still able to obtain smaller standard deviation than the standard conformal prediction which shows the adaptiveness of Algorithm 3 even without any knowledge of the domains. From Figure 3 (Right), we see that Algorithm 3 is able to maintain coverage, while ensuring low standard deviations across different $1-\alpha$.
@@ 6.3 Results for language task
Figure 4: Distribution of test recall. LlaMA-3-8B was calibrated with 3 different score functions and test data were labeled according to 4.2. Recall was calculated with the standard deviation plotted. The standard deviation is across 100 different test environments, obtained by sampling Dirichlet distribution with $\alpha^{\prime}=0.5$. Standard LLM uncertainty estimation method (blue) is sensitive to distribution shifts as evidenced by the high variance in recall across test-environments, while the recall with our method A3 (orange) tightly follows the desired target recall.
@@ 7 Conclusion, Limitations, and Future Work
8 Broader Impact
As people become increasingly dependent on AI models, uncertainty quantification for these models are ever more crucial. In most real-world settings, distribution shifts are almost impossible to avoid which necessitates uncertainty quantification methods that are adaptive to remedy the impacts. Our work examines the setting of test time distribution shifts which has a broad range of applications in deploying safe and trustworthy AI models.
Acknowledgments and Disclosure of Funding
YB and SPK were partially funded by Captil One CREDIF fellowship and grant respectively. SPK also thanks Yaodong Yu for initial conversations on this research.
@@ Appendix A Proofs
This proof follows the proof for Theorem 4.3 from [20] with some modifications. Suppose $\mathcal{E}$ is the event
=$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
Since $S(X^{k}_{1},Y^{k}_{1};\hat{f}),\cdots,S(X^{k}_{n_{k}},Y^{k}_{n_{k}};\hat{f}),S(X_{\text{test}},Y_{\text{test}};\hat{f})$ are exchangeable, we have that the above expression is lower bounded by
@@
A.2 Proof of Theorem 3.3
This proof follows the proof for Theorem 4.3 from [20] with some modifications. Suppose $\mathcal{E}$ is the event
=$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
Since $S(X^{k}_{1},Y^{k}_{1};\hat{f}),\cdots,S(X^{k}_{n_{k}},Y^{k}_{n_{k}};\hat{f}),S(X_{\text{test}},Y_{\text{test}};\hat{f})$ are exchangeable, we have that the above expression is lower bounded by
@@
A.3 Proof of Theorem 3.5
This proof follows the proof for Theorem 4.3 from [20] with some modifications. Suppose $\mathcal{E}$ is the event
=$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
Since $S(X^{k}_{1},Y^{k}_{1};\hat{f}),\cdots,S(X^{k}_{n_{k}},Y^{k}_{n_{k}};\hat{f}),S(X_{\text{test}},Y_{\text{test}};\hat{f})$ are exchangeable, we have that the above expression is lower bounded by
@@
Appendix B Overview of Score Functions
Although conformal prediction algorithms provide marginal coverage guarantees for arbitrary score functions, a poorly designed score function can lead to uninformative prediction sets. For our experiments we explore 3 different score functions for both vision and language tasks.
B.1 Vision Tasks
@@
  • Least Ambiguous Set-valued Classifier (LAC) [28]. Given data $(X,y)$ where $y$ is the true label of $X$, define $S(X,y;\hat{f})$ as
@@
  • Adaptive Prediction Set (APS) [26]. Given data $(X,y)$ where $y$ is the true label of $X$, define $S(X,y;\hat{f})$ as
@@
  • Regularized Adaptive Prediction Set (RAPS) [2]. Given data $(X,y)$ where $y$ is the true label of $X$, define $S(X,y;\hat{f})$ as
@@ B.2 Language Tasks
  • Length Normalized Scoring (LNS) [21]. Given a query $X$ and the generated response $\textbf{y}=\{y_{1},y_{2},\cdots,y_{L}\}$ of length $L$, define $S(X,\textbf{y};\hat{f})$ as the average log probability of the generated sequence, i.e.,
@@
  • Meaning-Aware Response Scoring (MARS) [5]. Given a query $X$ and the generated response $\textbf{y}=\{y_{1},y_{2},\cdots,y_{L}\}$ of length $L$, define $S(X,\textbf{y};\hat{f})$ as
@@
  • Degree Matrix Uncertainty [19]. We adopt the uncertainty estimate definition of [19] where the score only depends on the query $X$. Given a query $X$ and $m$ generated responses $\textbf{y}_{1},\textbf{y}_{2},\cdots,\textbf{y}_{m}$, first, define $W$ as a matrix of pairwise entailment dependencies where $W_{ij}$ represents the entailment dependency between output response $\textbf{y}_{i}$ and $\textbf{y}_{j}$. Entailment dependencies are calculated by using a Natural Language Inference classifier [13] that classifies generated responses into three classes: entailment, neutral, or contradiction. We then define the degree matrix $D$ as
@@
Appendix C Additional Experiments on Adapting to Distribution Shifts with Domain Knowledge
@@
Appendix D Additional Experiments on Adapting to Distribution Shifts without Domain Knowledge
@@
Appendix E Compute Resources

### v2

@@ 6.2 Results for vision tasks
Coverage with varying test environments. We obtain the results for Algorithm 3 with the same setup as section 5.2 and the results are plotted in Figure  3. From Figure  3 (Left) we observe that Algorithm 3 was able to provide coverage for the majority of test environments while standard conformal prediction could not for a significant number of test environments. Although not as small as Algorithm 1 and 2, Algorithm 3 is still able to obtain smaller standard deviation than the standard conformal prediction which shows the adaptiveness of Algorithm 3 even without any knowledge of the domains. From Figure 3 (Right), we see that Algorithm 3 is able to maintain coverage, while ensuring low standard deviations across different $1-\alpha$.
@@ 6.3 Results for language task
Figure 4: Controlling LLM hallucinations. LlaMA-3-8B was calibrated with 3 different score functions and test data were labeled according to 4.2. Recall was calculated with the standard deviation plotted. The standard deviation is across 100 different test environments, obtained by sampling Dirichlet distribution with $\alpha^{\prime}=0.5$. Standard LLM uncertainty estimation method (blue) is sensitive to distribution shifts as evidenced by the high variance in recall across test-environments, while the recall with our method A3 (orange) tightly follows the desired target recall.
@@ Appendix A Proofs
Proof.
Suppose we have the case of $k=2$ such that
$S(X_{\text{test}})\in\begin{cases}[0,1)&\text{if }X_{\text{test}}\sim\mathbb{P}_{1}\\ [1,2]&\text{if }X_{\text{test}}\sim\mathbb{P}_{2}.\end{cases}$ (4)
Furthermore, suppose $q_{\alpha,1}$ and $q_{\alpha,2}$ are the $1-\alpha$ quantile of scores from domain 1 and 2 respectively. In the extreme case that all $1-\gamma$ fraction of the mistakes that the domain classifier makes are on $X$’s such that $S(X)<q_{\alpha,2}$. It must be that these inputs are mis-covered because $S(X)>q_{\alpha,1}$. Therefore, the fraction of X that are covered is at most $1-\alpha-(1-\gamma)=\gamma-\alpha$. ∎
A.2 Proof of Theorem 3.1
This proof follows the proof for Theorem 4.3 from Lu et al. (2023) with some modifications. Suppose $\mathcal{E}$ is the event
$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
[…]
@@
A.3 Proof of Theorem 3.3
Suppose $\mathcal{E}$ is the event
$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
[…]
@@
A.4 Proof of Theorem 3.5
Suppose $\mathcal{E}$ is the event
$\mathcal{E}=\{\forall k\in[K],\exists\pi_{k},(S^{k}_{\pi_{k}(1)},\cdots S^{k}_{\pi_{k}(n_{k})},S^{k}_{\pi_{k}(n_{k}+1)})=(s^{k}_{1},\cdots,s^{k}_{n_{k}},s^{k}_{n_{k}+1})\},$
@@
[…]
@@
Appendix B Problem Setup
B.1 Background
Preliminaries.
We let $\mathcal{X}$ and $\mathcal{Y}$ denote the input and target space of a multiclass classification task. $\hat{f}:\mathcal{X}\rightarrow\Delta^{J}$ is the pre-trained classifier for the classification task and the output of $\hat{f}$ is a probability distribution over $J$ possible outcomes, e.g., the softmax output of a neural network. $\hat{f}(X)_{i}$ represents the $i$-th entry of the output of $\hat{f}$. We let $z:\mathcal{X}\rightarrow\mathbb{R}^{d}$ denote an embedding function that maps the input into a d-dimensional embedding space, and $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ stands for some similarity measure between embeddings.
We denote by $\mathbb{P}_{k}$ the distribution of the $k$-th domain, where there are $K$ domains in total. The calibration dataset from the $k$-th domain is represented as $\{(X^{k}_{i},Y^{k}_{i})\}^{n_{k}}_{i=1}$, and each pair $(X^{k}_{i},Y^{k}_{i})$ is assumed to be drawn i.i.d. from $\mathbb{P}_{k}$. The overall calibration dataset is sampled from the training environment $\mathbb{P}_{\text{train}}$, which is a mixture of the $K$ domain distributions. The score function is represented by $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ and will make use of $\hat{f}$. We define $m_{k}(q)$ as $|\{S(X^{k}_{i},Y^{k}_{i};\hat{f})\leq q\}|$, which is the number of number of calibration data in domain $k$ with score less than or equal to $q$.
Lastly, the test data is denoted by $(X_{\text{test}},Y_{\text{test}})$, drawn from a test environment $\mathbb{P}_{\text{test}}$, which is an unknown mixture of the $k$ domains $\mathbb{P}_{k}$. We denote the set of all possible test environments by $\mathcal{D}$. Examples of such test environments are illustrated in Figure 1.
Conformal Prediction Under Exchangeability
Given a test data $X_{\text{test}}$ with unknown label $Y_{\text{test}}$, a calibration set $\{(X_{i},Y_{i})\}^{n}_{i=1}$, which is distinct from train and test set, and a user defined error rate $\alpha$, the goal of conformal prediction is to build a prediction set $C_{\alpha}$ that satisfies 1. To conformalize a model to output a valid prediction set, the following procedure is followed: First, a score function $S$ is defined. Second, the threshold $\hat{q}_{\alpha}$ is computed as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of $\{S(X_{i},Y_{i};\hat{f})\}_{i=1}^{n}$. Lastly, the prediction set $C_{\alpha}(X_{test})$ is returned such that $C_{\alpha}(X_{test})=\{y|S(X_{test},y;\hat{f})\leq\hat{q}_{\alpha}\}$. If the calibration data and the test data are drawn i.i.d from the some domain, then $C_{\alpha}(X_{test})$ satisfies the marginal coverage guarantee due to exchangeability between calibration and test data (Angelopoulos & Bates, 2021). We refer to this method as the standard or unweighted conformal prediction throughout the paper.
Conformal Risk Control
The conformal prediction framework can be extended to provide guarantee beyond coverage. Given a prediction set $C(X_{\text{test}})$, a loss function $\ell$ that decreases as $|C(X_{\text{test}})|$ increases, and a user defined error rate $\alpha$, the conformal risk control guarantee is defined as,
$\mathbb{E}[\ell(C(X_{\text{test}}),Y_{\text{test}})]\leq\alpha$ (8)
(Angelopoulos et al., 2023). Note that the marginal coverage guarantee can be reduced to 8 if we define $\ell$ as the miscoverage loss, i.e., $\ell(C(X_{\text{test}}),Y_{\text{test}})=\mathbf{1}\{Y_{\text{test}}\notin C(X_{\text{test}})\}$. We refer to Angelopoulos et al. (2023) for the details of conformal risk control. One application for the conformal risk control framework is in large language model (LLM) uncertainty estimation, in particular, hallucination detection. Hallucination refers to when an LLM generate responses that are factually false or inconsistent with the training data. Conformal risk control can be used to select a threshold to determine whether an LLM output is a hallucination or not while maintaining a theoretical bound to metrics such as sensitivity or precision.
Appendix C Overview of Score Functions
A (conformal) score function maps an input pair $(X,Y)$ to a real-valued score. A larger score indicates less conformity between $(X,Y)$ and other training data. Although conformal prediction algorithms provide marginal coverage guarantees for arbitrary score functions, a poorly designed score function can lead to uninformative prediction sets. For our experiments we explore 3 different score functions for both vision and language tasks.
C.1 Vision Tasks
@@
[…]
@@
[…]
@@
[…]
@@ C.2 Language Tasks
[…]
@@
[…]
@@
  • Degree Matrix Uncertainty (Lin et al., 2024). We adopt the uncertainty estimate definition of Lin et al. (2024) where the score only depends on the query $X$. Given a query $X$ and $m$ generated responses $\textbf{y}_{1},\textbf{y}_{2},\cdots,\textbf{y}_{m}$, first, define $W$ as a matrix of pairwise entailment dependencies where $W_{ij}$ represents the entailment dependency between output response $\textbf{y}_{i}$ and $\textbf{y}_{j}$. Entailment dependencies are calculated by using a Natural Language Inference classifier (He et al., 2021) that classifies generated responses into three classes: entailment, neutral, or contradiction. We then define the degree matrix $D$ as
@@
Appendix D Additional Experiments on Adapting to Distribution Shifts with Domain Knowledge
@@
Appendix E Additional Experiments on Adapting to Distribution Shifts without Domain Knowledge
@@
Appendix F Compute Resources